the length of a rectangular garden is 4m more than 3 times its width. If the perimeter of the garden is 56m, what are the dimesions of the garden?
Answer by mananth(16946)

the length of a rectangular garden is 4m more than 3 times its width. If the perimeter of the garden is 56m, what are the dimesions of the garden?
...
let width be x
length = 3x+4
..
2(L+W)=perimeter
2(x+3x+4)=56
2(4x+4)=56
/2
4x+4=28
4x=28-4
4x=24
/4
x=6m the with
Length = 3x+4 = 3*6+4 = 22m
..
m.ananth@hotmail.ca